OpenStudy (zab505):

Solve equation 2/3x-1/2=1/6-2x Medal & Fan.

OpenStudy (zab505):

Solve for x

OpenStudy (nurali):

2/3x+2x=1/6+1/2 2/3x+2x=2/3 x(2/3+2)=2/3 8/3x=2/3 x=2/3*3/8 x=2/8 x=1/4

OpenStudy (zab505):

Thank you so much!

OpenStudy (nurali):

Anytime.

OpenStudy (zab505):